# OC3D: Weakly Supervised Outdoor 3D Object Detection with Only Coarse Click Annotation

### v2

@@ Abstract
LiDAR-based outdoor 3D object detection has received widespread attention. However, training 3D detectors from the LiDAR point cloud typically relies on expensive bounding box annotations. This paper presents OC3D, an innovative weakly supervised method requiring only coarse clicks on the bird’s eye view of the 3D point cloud. A key challenge here is the absence of complete geometric descriptions of the target objects from such simple click annotations. To address this problem, our proposed OC3D adopts a two-stage strategy. In the first stage, we initially design a novel dynamic and static classification strategy and then propose the Click2Box and Click2Mask modules to generate box-level and mask-level pseudo-labels for static and dynamic instances, respectively. In the second stage, we design a Mask2Box module, leveraging the learning capabilities of neural networks to update mask-level pseudo-labels, which contain less information, to box-level pseudo-labels. Experimental results on the widely used KITTI and nuScenes datasets demonstrate that our OC3D with only coarse clicks achieves state-of-the-art performance compared to weakly-supervised 3D detection methods. Combining OC3D with a missing click mining strategy, we propose an OC3D++ pipeline, which requires only 0.2% annotation cost in the KITTI dataset to achieve performance comparable to fully supervised methods. The code will be made publicly available.
Introduction
@@
In recent years, notable progress has been made in LiDAR-based 3D object detection research (Wu et al. 2023; Huang et al. 2024). Despite these advancements, the need for precise bounding box supervision remains a major challenge due to its time-consuming and labor-intensive nature. For instance, the KITTI dataset (Geiger, Lenz, and Urtasun 2012) contains 3,712 training scenes with over 15,000 vehicle instances, where manual annotation of a single instance can take roughly 114 seconds (Meng et al. 2021). The extensive labeling efforts required escalate dramatically when scaling detectors to larger-scale datasets (Sun et al. 2020; Caesar et al. 2020; Mao et al. 2021), thereby hindering further research in a fully supervised manner.
Figure 1: (a) 6DoF box annotation; (b) Unlike costly box annotations, coarse-click alternatives require only quick clicks on objects in the 2D BEV plane, yet offer limited supervision; (c) An empirical study on maximizing the utility of weak click labels reveals significant performance gains when separating static and dynamic objects (yellow bars) compared to directly estimating 3D bounding boxes from clicks (blue bars) or discarding pseudo labels for dynamic objects (orange bars). The proposed OC3D (green bars) further refines label quality via Mask2Box conversion, substantially enhancing detection performance by 17%.
@@
To alleviate the annotation burden, recent studies have explored alternatives that require fewer annotated frames or instances to train high-performing 3D object detectors. Specifically, semi-supervised methods (Wang et al. 2021; Park et al. 2022; Liu et al. 2023a) leverage a subset of the annotated frames, while sparsely-supervised methods (Xia et al. 2023b; Liu et al. 2022a; Xia et al. 2024) rely on only one bounding box annotation per frame during training. While these approaches have significantly lowered annotation costs, annotating 6DoF bounding boxes for every scene remains time-consuming.
@@
To provide a faster, albeit less precise, method of human supervision, WS3D (Meng et al. 2021) and ViT-WSS3D (Zhang et al. 2023) propose a mixed supervision strategy that replaces some box annotations with the center-click annotations. In this approach, annotators click the center of objects on the Bird’s Eye View (BEV) to generate center-level labels, reducing labeling time per instance to approximately 2.5 seconds, which is 50 times faster than traditional bounding box labeling.
@@
However, relying on a single center point has significant limitations: (1) it requires annotators to precisely indicate the center position and (2) it fails to accurately represent the shape and scale of objects, especially in sparse point clouds where objects may be partially observed. These challenges become even more severe for moving objects, where motion across frames further complicates the estimation of an accurate bounding box from clicks. As a result, prior works have struggled to scale up the use of click supervision, combined with traditional box annotation for mixed supervision, which diminishes the overall effectiveness.
In this paper, we introduce a novel pure click-supervised approach for 3D object detection (OC3D), which employs temporal cues to distinguish between static and moving instances and progressively recover their box supervision from coarse clicks (Fig. 1 (b)) for detector training. This strategy offers greater flexibility than center-clicked approaches by accommodating inaccurate clicks through shift tolerance, which can be corrected via multiple-frame consistency. Our framework comprises three key designs: (1) To efficiently classify the motion state of provided clicks, we analyze the point density at the clicked location from a statistical perspective, noting that vary point density over a short period for moving objects while stable density for static instances. (2) For static objects, we implement a Click2Box strategy that aggregates the neighboring points across multiple frames to reconstruct the object’s structure and lift click-level supervision to a precise 3D box. (3) For moving objects, which are more challenging to regress, we generate mask-level pseudo labels (Click2Mask) to assign the clustered points with labels, enabling models to learn to predict the object’s location. To compensate for the lack of supervision in predicting shapes and scales for moving objects, we leverage their geometric similarity as a bridge, further refining the 3D box through a Mask2Box process. To address extreme cases where only one click is provided per scene (sparsely supervised setting), we propose a variant OC3D++, which identifies missed click instances and facilitates training with only 0.2% of the annotations required for full labeling. We evaluated OC3D and OC3D++ on the widely adopted KITTI and nuScenes datasets. Remarkably, OC3D achieves competitive performance with weakly-supervised baselines that rely on accurate box annotations. Furthermore, when the number of clicks is reduced to one per scene, OC3D++ still maintains robust performance, demonstrating the resilience and effectiveness of our approach under sparely supervised conditions.
In summary, our contributions are:
@@
  • We propose the first method of only click annotated 3D object detection from point cloud (OC3D), which solely relies on coarse clicks on the BEV maps. This approach dramatically reduces the annotation cost of 3D object detection tasks to 1% $\sim$ 0.2%.
  • We design the Click2Box and Click2Mask modules according to the motion attributes of instances, inferring accurate mixed supervision information from the click annotations.
  • We design a Mask2Box module that upgrades the less informative mask-level pseudo labels to box-level pseudo labels, which compensates for the loss of object shape and scale information in mask-level supervision.
Related Work
@@
In recent years, fully-supervised 3D object detection has been widely studied. The early methods (Lang et al. 2019; Yan, Mao, and Li 2018; Yin, Zhou, and Krähenbühl 2021) utilized an end-to-end one-stage object detection strategy, predicting detection boxes directly from point clouds. Subsequently, the two-stage methods (Shi, Wang, and Li 2019; Deng et al. 2021; Shi et al. 2023; Wu et al. 2022b; Chen et al. 2023; Wang et al. 2024) introduced an additional proposal box refinement stage, which improves detection performance by refining regions of interest. Meanwhile, multi-stage methods (Wu et al. 2022a; Cai et al. 2022) progressively refined proposals using a cascaded optimization strategy, leading to more precise predictions. Despite achieving excellent performance, all these methods require costly box annotations, the generation of which is time-consuming and labor-intensive.
Figure 2: The pipeline of the proposed OC3D. (a) Initially, a novel motion state classification strategy is introduced, followed by the generation of mask-level pseudo-label $\mathbf{A}_{m}$ and box-level pseudo-label $\mathbf{A}_{b}$, utilizing the Click2Mask and Click2Box modules, respectively. (b) With the mixed pseudo-labels generated by stage (a), train the detector and then update the mask-level supervision to box-level based on high-confidence predictions of the trained detector.
Label-efficient 3D Object Detection.
Research on reducing annotation costs in 3D object detection tasks has received widespread attention. Around semi-supervised methods (Liu et al. 2023a; Park et al. 2022; Wang et al. 2021), they selected only a small number of fully annotated frames as labeled data, using the remaining frames as unlabeled data. These methods used teacher-student networks for distillation learning to mine and generate pseudo-labels. Sparsely-supervised methods (Xia et al. 2023b; Liu et al. 2022a) adopted a sparsely annotated strategy, retaining only one complete bounding box label for each selected frame. They utilized specially designed unlabeled object mining modules to discover potential pseudo-labels. There are also some multimodal weakly supervised object detection methods (Wei et al. 2021; Liu et al. 2022b) that used 2D bounding boxes instead of 3D bounding boxes to carry out object detection tasks. Although these strategies have significantly reduced the dependence on 3D boxes, it is still not possible to completely abandon laborious box-level annotations.
@@
Click-level annotation is a very labor-saving strategy for labeling. Weakly supervised segmentation methods (Liu, Qi, and Fu 2021; Tao et al. 2022; Liu et al. 2023b) have explored the application of only click-level annotation in segmentation tasks. However, this annotation strategy for 3D object detection is still under-explored. WS3D (Meng et al. 2021) first introduced click-level annotation into weakly supervised 3D object detection, using partial box-level supervision to guide the inference of click-level annotation towards bounding box estimation. ViT-WSS3D (Zhang et al. 2023) leveraged vision transformer (Dosovitskiy et al. 2020) to further optimize this process. MixSup (Yang, Fan, and Zhang 2024) made three coarse clicks around an object to generate point cluster labels and proposes a mixed supervision training strategy by combining these cluster-level labels with box-level labels. Current weakly supervised object detection algorithms have successfully incorporated click-level annotations as a substitute for some box-level annotations, but they cannot achieve 3D object detection tasks using only click-level annotations.
Proposed Approach
@@
We start by defining the task of weakly-supervised 3D object detection using pure click supervision. In this setting, the detector is trained exclusively on instance-level click annotations $c_{o}=\left\{x_{o},y_{o}\right\}$ on the 2D BEV plane, and is then required to predict the full 3D bounding boxes $\left\{x,y,z,l,w,h,\theta\right\}$. Here, $x,y,z$ represent the coordinates of the 3D center point, while $l,w,h$ denote the object’s dimensions, and $\theta$ specifies its orientation. Importantly, the click annotations employed are not limited to precise center clicks; further details are provided in the supplementary material. As shown in Fig. 2, given the coarse nature of click supervision, our approach proceeds by first (1) generating mixed-granularity supervision for both static and moving objects using the Click2Box and Click2Mask strategies, followed by (2) applying Mask2Box-enhanced self-training to refine and improve the quality of generated supervision. The specific steps involved in this process are elaborated as follows.
Mixed Pseudo-Label Generation
To accurately estimate 3D bounding boxes from click annotations as pseudo labels, we aim to leverage temporal cues to enrich sparse point annotations by aggregating registered points across consecutive frames. For static instances, the aggregation of dense points facilitates the complete capture of spatial geometric information of the instance. However, for dynamic objects, the dense point aggregation often fails to capture the full geometric structure, leading to low-quality pseudo-labels. This challenge motivates us to first classify the motion status of instances corresponding to each annotation click:
@@
We observe the duration of local point distribution at clicked positions during a long sequence traversal. Specifically, for static instances, the local point cloud at the clicked position exhibits a continuous distribution, whereas, for dynamic instances, the local point cloud at the clicked position is transient throughout the traversal. Motivated by this, we utilize the persistence of points at local positions within the long sequence for dynamic and static classification.
For each click annotation $c_{o}=(x_{o},y_{o})$ at the $t$-th frame, we gather adjacent frames =$\mathcal{F}=\left\{f_{t-k},...,f_{t},...,f_{t+k}\right\}$ within a local time window $k$, followed by ground removal (Himmelsbach, Hundelshausen, and Wuensche 2010) as a preprocessing step. Our primary focus is on the BEV points in $\mathcal{F}$, denoted as $\{\mathbf{P}^{\operatorname{BEV}}_{t}\in\mathbb{R}^{N\times 2}\}_{t\in[t-k,t+k]}$, where $N$ indicates the number of BEV points in each frame. To determine the persistence of the clicked position $(x_{o},y_{o})$, we search for its neighboring BEV points within a radius $r$, resulting in the collection $\{\mathbf{N}_{t}\}_{t\in[t-k,t+k]}$, with each time step having a cardinality of $N_{t}$:
@@
For static objects, dense points express complete geometric structures, which supports the fitting of high-quality bounding box pseudo-labels from the point cloud distribution. Motivated by this observation, we concatenate the neighboring points of multiple frames $\{\mathbf{N}_{t}\}_{t\in[t-k,t+k]}$ to obtain local dense points $\mathbf{D}_{t}$ for the time step $t$. We perform DBSCAN (Ester et al. 1996) clustering algorithm on $\mathbf{D}_{t}$ to generate several discrete point clusters. We retain the cluster of points whose center is closest to the clicked position and consider the points in this cluster as the foreground points of the clicked instance. Finally, we perform a bounding box fitting algorithm (Zhang et al. 2017) on the foreground points to generate a box-level pseudo-label. We utilize Click2Box to infer box-level pseudo-labels $\mathbf{A}_{b}$ from the click annotations of all static instances.
@@
For dynamic objects, we opt to leverage only the single-frame point cloud $\mathbf{P}_{t}$ due to the long-tail distribution observed in aggregated points resulting from motion differences (Chen et al. 2022). Although the instance shape and scale cannot be revealed by click-level labels, the foreground points from click-annotated frame can still provide reliable semantic information and coarse location information. Consequently, instead of generating box-level pseudo labels for dynamic objects, we produce mask-level pseudolabels by extracting the foreground points in $\mathbf{P}_{t}$. To identify these foreground points, we employ the DBSCAN clustering algorithm on raw point clouds, and select the cluster with the center closest to the click location a mask-level pseudo label. The resulting mask-level pseudo labels denoted as $\mathbf{A}_{m}$, are derived from the click annotations of all dynamic instances.
Mask2Box Enhanced Mixed Supervised Training
In contrast to traditional 3D object detectors that are solely reliant on box supervision, our approach delves into weakly supervised detectors with mixed supervision. Inspired by MixSup(Yang, Fan, and Zhang 2024), we re-engineered the strategy for supervision allocation. However, the information about mask-level supervision is limited, and it is challenging to achieve optimal detector performance. To address this issue, combining with an iterative training, we introduce a mask2box enhanced training strategy, which leverages the high-confidence outputs from the last iterative detector to refine mask-level pseudo-labels into more accurate bounding box-level pseudo-labels, thereby improving the overall precision of the detection process.

### v3

# SC3D: Label-Efficient Outdoor 3D Object Detection via Single Click Annotation

@@ Abstract
LiDAR-based outdoor 3D object detection has received widespread attention. However, training 3D detectors from the LiDAR point cloud typically relies on expensive bounding box annotations. This paper presents SC3D, an innovative label-efficient method requiring only a single coarse click on the bird’s eye view of the 3D point cloud for each frame. A key challenge here is the absence of complete geometric descriptions of the target objects from such simple click annotations. To address this issue, our proposed SC3D adopts a progressive pipeline. Initially, we design a mixed pseudo-label generation module that expands limited click annotations into a mixture of bounding box and semantic mask supervision. Next, we propose a mix-supervised teacher model, enabling the detector to learn mixed supervision information. Finally, we introduce a mixed-supervised student network that leverages the teacher model’s generalization ability to learn unclicked instances. Experimental results on the widely used nuScenes and KITTI datasets demonstrate that our SC3D with only coarse clicks, which requires only 0.2% annotation cost, achieves state-of-the-art performance compared to weakly-supervised 3D detection methods. The code will be made publicly available.
1 Introduction
Figure 1: (a) Unlike traditional costly box annotations, coarse-click alternatives require only quick clicks on single object in the 2D BEV plane, yet offer limited supervision; (b) demonstrates a comparison with existing label-efficient methods on the KITTI dataset, where we reduce the annotation cost to 0.2%, enhancing the previous scheme’s labeling efficiency by over 10 times, while still maintaining the performance of the detector.
@@ 1 Introduction
In recent years, notable progress has been made in LiDAR-based 3D object detection research [39, 13, 4, 34]. Despite these advancements, the need for precise bounding box supervision remains a major challenge due to its time-consuming and labor-intensive nature. For instance, the KITTI dataset [10] contains 3,712 training scenes with over 15,000 vehicle instances, where manual annotation of a single instance can take roughly 114 seconds [24]. The extensive labeling efforts required escalate dramatically when scaling detectors to larger-scale datasets [31, 1, 23], thereby hindering further research in a fully supervised manner.
@@
To alleviate the annotation burden, recent studies have explored alternatives that require fewer annotated frames or instances to train high-performing 3D object detectors. Specifically, semi-supervised methods [33, 25, 18] leverage a subset of the annotated frames, while sparsely-supervised methods [42, 16, 43] rely on only one bounding box annotation per frame during training. While these approaches have significantly lowered annotation costs, annotating 6DoF bounding boxes for every scene remains time-consuming.
@@
To provide a faster, albeit less precise, method of human supervision, WS3D [24] and ViT-WSS3D [50] propose a mixed supervision strategy that replaces some box annotations with the center-click annotations. In this approach, annotators click the center of objects on the Bird’s Eye View (BEV) to generate center-level labels, reducing labeling time per instance to approximately 2.5 seconds, which is 50 times faster than traditional bounding box labeling.
@@
However, relying on full coverage of the central point has significant limitations: (1) it requires annotators to precisely indicate the center position for all instance in each frame and (2) it fails to accurately represent the shape and scale of objects, especially in sparse point clouds where objects may be partially observed. These challenges become even more severe for moving objects, where motion across frames further complicates the estimation of an accurate bounding box from clicks. As a result, prior works have struggled to scale up the use of sparse click supervision, combined with traditional box annotation for mixed supervision, which diminishes the overall effectiveness.
In this paper, we introduce a label-efficient 3D object detection approach, SC3D, which performs only sparse click annotations (Fig. 1), greatly reducing the annotation cost. Our framework comprises three key designs: (1) To recover accurate supervisions from coarse click annotations, we use temporal cues to determine the status of the clicked instances and then generate mixed-supervisions. (2) To optimize learning from mixed supervision, we designed a mixed-supervised teacher network to capture the fundamental patterns within the mixed supervision data. (3) To learn from unclicked instances, we introduce a mixed-supervised student network that uses the teacher’s generalization to capture their patterns. We evaluated SC3D on the widely adopted nuScenes and KITTI datasets. Remarkably, SC3D achieves competitive performance with weakly-supervised baselines that rely on accurate box annotations, demonstrating the effectiveness of our approach under sparely click conditions. In summary, our contributions are:
@@
  • We propose the first method of sparse click annotated outdoor 3D object detection (SC3D), which solely relies on a single coarse click on each frame. This approach dramatically reduces the annotation cost of 3D object detection tasks to 0.2%.
  • We design a mixed pseudo-label generation module that combines temporal cues and local point cloud distribution to recover accurate supervision from coarse click annotations.
  • We propose a mixed-supervised teacher-student network that expands the information of unclicked instances from limited mixed supervision, thereby enhancing the performance of the detector.
2 Related Work
@@
In recent years, fully-supervised 3D object detection has been widely studied. The early methods [15, 44, 47, 53, 11] utilized an end-to-end one-stage object detection strategy, predicting detection boxes directly from point clouds. The advantage of these approaches is that the detector inference speed is fast, which more easily meets the real-time requirements of the autonomous driving. To improve the performance of 3D detector, the two-stage methods [27, 5, 29, 38, 4, 35, 21, 34] introduced an additional proposal refinement stage, which improves detection performance by refining regions of interest. Meanwhile, multi-stage methods [37, 2] progressively refine detection proposals through a series of stages, each employing a cascaded optimization strategy. This iterative process leads to increasingly accurate predictions as the algorithm refines the initial detections, ultimately resulting in more precise localization and classification of 3D objects within the scene. Despite achieving excellent performance, all these methods require costly box annotations, the generation of which is time-consuming and labor-intensive.
3D Object Annotator.
Research on reducing annotation costs in 3D object detection tasks has received widespread attention. To achieve optimal automatic annotation for 3D instances, offboard 3D object detectors [8, 26, 22, 45, 49] utilize manual annotations and future frame information to train a robust annotator and bounding box corrector. However, the success of offboard detectors still relies heavily on a large amount of manual annotation. To reduce dependence, the semi-supervised methods [18, 25, 33] select only a small number of fully annotated frames as labeled data, using the remaining frames as unlabeled data. These methods used teacher-student networks for distillation learning to mine and generate pseudo-labels. In addition, sparsely-supervised methods employ an alternative approach to reduce dependency. These methods [42, 16, 43] adopted a sparsely annotated strategy, retaining only one complete bounding box label for each selected frame. They utilized specially designed unlabeled object mining modules to discover potential pseudo-labels. There are also some multimodal weakly supervised object detection methods [36, 17, 9] that used 2D bounding boxes instead of 3D bounding boxes to carry out object detection tasks. Although these strategies have significantly reduced the dependence on 3D boxes, it is still not possible to completely abandon laborious box-level annotations.
Figure 2: The overview of proposed SC3D. (a) Initially, a novel motion state classification strategy is introduced, followed by the generation of box-level pseudo-label ${\color[rgb]{1,.5,0}\mathbf{L}_{b}}$ and mask-level pseudo-label ${\color[rgb]{0.546875,0,0}\mathbf{L}_{m}}$, utilizing the Click2Mask and Click2Box modules, respectively. (b) With the mixed pseudo-labels generated by stage (a), train the mixed-supervised teacher detector and then update the mask-level supervision to box-level based on high-confidence predictions. (c) Utilizing the generalization of the teacher network to produce mixed pseudo-labels for unlabeled instances, further enhancing the performance of the mixed-supervised student network.
@@
Click-level annotation is a very labor-saving strategy for manual labeling. Weakly supervised segmentation methods [20, 32, 19] have explored the application of only click-level annotation in segmentation tasks. However, this annotation strategy for 3D object detection is still under-explored. WS3D [24] first introduced click-level annotation into weakly supervised 3D object detection, using partial box-level supervision to guide the inference of click-level annotation towards bounding box estimation. Based on this, ViT-WSS3D [50] leveraged vision transformer [6] to further optimize this process. Mixsup [46] made three coarse clicks around an object to generate point cluster labels and proposes a mixed supervision training strategy by combining these cluster-level labels with box-level labels. Current weakly supervised object detection algorithms have successfully incorporated click-level annotations as a substitute for some box-level annotations, but they cannot achieve 3D object detection tasks using only click-level annotations. To address this issue, we design a mixed pseudo-label generation module to efficiently convert click annotations into supervision.
3 Proposed Approach
@@
We start by defining the task of weakly-supervised 3D object detection using sparse click supervision. In this setting, the detector is trained exclusively on a single instance-level click annotation $c_{o}=\left\{x_{o},y_{o}\right\}$ on each frame of 2D BEV plane, and is then required to predict the full 3D bounding boxes $\left\{x,y,z,l,w,h,\theta\right\}$. Here, $x,y,z$ represent the coordinates of the 3D center point, while $l,w,h$ denote the object’s dimensions, and $\theta$ specifies its orientation. Importantly, the click annotations employed are not limited to precise center clicks; further details are provided in the supplementary material. As shown in Fig. 2, (a) mixed pseduo-label generation; (b) mixed-supervised teacher training; and (c) mixed-supervised student training. The specific steps involved in this process are elaborated as follows.
3.1 Mixed Pseudo-Label Generation
To accurately estimate 3D bounding boxes from click annotations as pseudo labels, we aim to leverage temporal cues to enrich sparse point annotations by aggregating registered points across consecutive frames. For static instances, the aggregation of dense points facilitates the complete capture of spatial geometric information. However, for dynamic objects, the dense point aggregation often fails to capture the full geometric structure, leading to low-quality pseudo-labels. This challenge motivates us to first classify the motion status of instances corresponding to each click.
@@ 3.1 Mixed Pseudo-Label Generation
As shown in Fig. 3, we observe the duration of local point distribution at clicked positions during a long sequence traversal. Specifically, for static instances, the local points at the clicked position exhibits a continuous distribution, whereas, for dynamic instances, the local points is transient throughout the traversal. Motivated by this, we utilize the persistence of points at local positions within the long sequence for dynamic and static classification.
For each click annotation $c_{o}=(x_{o},y_{o})$ at the $t$-th frame, we gather adjacent frames =$\mathcal{F}=\left\{f_{t-k},...,f_{t},...,f_{t+k}\right\}$ within a local time window $k$, followed by ground removal [12]. Our primary focus is on the BEV points in $\mathcal{F}$, denoted as $\{\mathbf{P}^{\operatorname{BEV}}_{t}\in\mathbb{R}^{N\times 2}\}_{t\in[t-k,t+k]}$, where $N$ indicates the number of BEV points in each frame. To determine the persistence of the clicked position $(x_{o},y_{o})$, we search for its neighboring BEV points within a radius $r$, resulting in the collection $\{\mathbf{N}_{t}\}_{t\in[t-k,t+k]}$, with each time step having a cardinality of $N_{t}$:
@@
Figure 3: The difference in point distribution between dynamic and static objects in consecutive frames are as follows: dynamic objects have a rapid change in local point density over time; static objects have a stable local point density.
@@
For static objects, motivated by [40], dense points express complete geometric structures, which supports the fitting of high-quality bounding box pseudo-labels from the point cloud distribution. Therefore, we concatenate the neighboring points of multiple frames $\{\mathbf{N}_{t}\}_{t\in[t-k,t+k]}$ to obtain local dense points $\mathbf{D}_{t}$ for the time step $t$. We perform DBSCAN [7] clustering algorithm on $\mathbf{D}_{t}$ to generate several discrete point clusters. We retain the cluster of points whose center is closest to the clicked position and consider the points in this cluster as the foreground points of the clicked instance. Finally, we perform a bounding box fitting algorithm [52] on the foreground points to generate a box-level pseudo-label. We utilize Click2Box to infer box-level pseudo-labels $\mathbf{L}_{b}$ from the click annotations of all static instances.
@@
For dynamic objects, we opt to leverage only the single-frame point cloud $\mathbf{P}_{t}$ due to the long-tail distribution observed in aggregated points resulting from motion differences [3]. Although the instance shape and scale cannot be revealed by click-level labels, the foreground points from click-annotated frame can still provide reliable semantic information and coarse location information. Consequently, instead of generating box-level pseudo labels for dynamic objects, we produce mask-level pseudolabels by extracting the foreground points in $\mathbf{P}_{t}$. To identify these foreground points, we employ the DBSCAN clustering algorithm on raw point clouds, and select the cluster with the center closest to the click location a mask-level pseudo label. The resulting mask-level pseudo labels denoted as $\mathbf{L}_{m}$, are derived from the click annotations of all dynamic instances.
3.2 Mixed-Supervised Teacher Training
In contrast to traditional 3D object detectors that are solely reliant on box supervision, our approach delves into weakly supervised detectors with mixed supervision. Inspired by MixSup[46], we re-engineered the strategy for supervision allocation. However, the information about mask-level supervision is limited, and it is challenging to achieve optimal detector performance. To address this issue, combining with an iterative training, we introduce a mask2box enhanced teacher training strategy, which leverages the high-confidence outputs from the last iterative detector to refine mask-level pseudo-labels into more accurate bounding box-level pseudo-labels, thereby improving the overall precision of the teacher detector.